# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

### 11 Feb 2024 v1

@@ 2 Related Work
COC has been referred to as Legal Judgment Prediction (LJP) in previous research Medvedeva et al. (2020). Under both terms, there exist numerous works involving decision corpora from the European Court of Human Rights (ECtHR) Aletras et al. (2016); Chalkidis et al. (2019); Tyss et al. (2023). All of these approaches mainly focus on the COC performance of models, which is commonly measured against the human majority class. Our SV-ECtHR dataset extends this work and contains the nuanced judges’ vote information. We proceed to use it to systematically investigate disagreement among judges and the alignment of perceived difficulty between models and judges.
@@
Calibration Existing calibration studies primarily focus on a classifier’s awareness of its predictive performance, commonly evaluated against the human majority class Guo et al. (2017); Desai and Durrett (2020). Baan et al. 2022 argue that calibrating against the human majority class is not meaningful in settings with inherent human disagreement. In our work, we assess the performance and human calibration of COC models in the context of split-vote, a naturally occurring human disagreement. To the best of our knowledge, this is the first systematic exploration of human calibration in the legal domain.
@@ 3 Legal and Linguistic Backgrounds
The Legal Decision Making Process at the ECtHR begins with the applicants lodging their accusation, alleging one or more violations of articles of the European Convention on Human Rights (ECHR). This allegation process corresponds to Task B (allegation prediction) in the widely used LexGLUE benchmark Chalkidis et al. (2022a). After receiving the complaint, the court undertakes a review of the case, aiming to determine whether a violation has indeed occurred, corresponding to Task A (violation prediction) in LexGLUE. Cases falling under the purview of well-established ECtHR case-law are directed to a three-judge Committee, while others may find themselves before a seven-judge Chamber, where decisions are reached through a majority vote. In some circumstances, the Grand Chamber, comprising 17 judges, adjudicates cases referred to it by request. In the process of rendering a judgment, the Court typically examines only the specific articles alleged by the applicant. This procedure aligns with Task A|B (violation prediction given allegation), as introduced by Santosh et al. 2022. Consequently, our study evaluates the uncertainty of COC models for Task A|B, mirroring the real legal process. Sources of Disagreement in Legal Decisions Recent research has delved into sources of HLV and disagreement across various NLP tasks from a linguistic perspective. The classic framework “Triangle of Reference" Ogden and Richards (1923) has been widely adopted to categorize disagreement sources in classification tasks Aroyo and Welty (2015); Jiang and Marneffe (2022). This concept primarily addresses the relationship between linguistic symbols and the corresponding objects they represent. To study uncertainty sources in NLG tasks, Baan et al. 2023 extend it to the “Double Triangle of Language Production” catering to the complexities of language generation. However, the legal decision-making process introduces an additional layer of complexity. Legal scholars conceptualize the decision-making process in case-law as circular Ichim (2019). To better capture this nature, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4 in App A), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the continuous reproduction of the main normative model over time.
Figure 2: Taxonomy of disagreement sources among judges in legal decisions and proxy labels.
@@ 4 The SV-ECHR Dataset
Dataset Analysis: We then merge the vote distribution to the ECHR A|B dataset Santosh et al. (2022) according to the document ID. We name the merged dataset SV-ECHR. It consists of 11k case fact descriptions along with target label information about which convention articles have been alleged to be violated (task B), and which the court has eventually found to be violated (task A), and the judges vote distribution of each alleged article. The dataset is chronologically split into training (2001–2016), validation (2016–2017) and test set (2017-2019) with 9k, 1k, 1k cases respectively. The label set includes 10 prominent ECHR convention articles. On average, each case has around 1.6 alleged articles. Among all 17,604 alleged case-alleged article pairs (hereafter pair), only round 7% are split-voted. These numbers underscore the significant label imbalance within the dataset. Additional statistics can be found in Tab 5 in App D. To investigate the extent of judges’ disagreement in SV-ECHR, we assess the entropies of vote distributions for each pair. For the detailed calculation and histogram of the entropy distribution, see App F. We see a large share ($\sim$ 60% of chamber votes) of single dissenting votes, where only one single judge voted differently than the six judge majority. A similar pattern is observed for Grand Chamber cases involving 17 judges. This pattern aligns with other observations made in empirical legal scholarship (Fobbe 2022, see App G).
@@ 5.1.1 The Data
Genuine Ambiguity is attributed to Normative Uncertainty of case law in the context of legal NLP (Xu et al. 2023b), which emerges when the court is presented with the possibility of justifying an outcome through multiple legal source interpretations and argumentation. Its occurrence is not uncommon in ECtHR judgments due to the deliberate drafting of the convention in a ‘flexible’ manner to allow tailoring the interpretation to domestic specificities based on the subsidiarity and margin of appreciation principles. It should be noted that there exists and analog factual ambiguity where the facts of the case are unclear based on limited (or contradicting) evidence provided. As the ECtHR does not engage in evidentiary reasoning, this aspect is out of scope for this work.
@@ 5.1.2 The Annotator/Judge
A comparable liberal vs conservative investigation in the ECtHR context would not be useful. It is composed of 46 Member States, each with its own spectrum of political ideologies, more pluralist and complicated than the liberal-conservative divide. We account for some such disagreement based on the democratic score of the State they are representing (the HighRep variable, see below). Moreover, there is no temporal stability in the judges’ behavior because they have a limited mandate and act in different bench formations (chamber, grand chamber, etc.). Focusing on individual ideologies to the exclusion of variables accounting for the wider court dynamics could not support accurate analysis. We have therefore chosen to focus instead on normative uncertainty and the specific legal context.
@@ 5.3 Proxies & Disagreement Correlation
HighRepCountry (High Reputation Country): Previous legal and political science work indicates ECtHR Judges tend to have more split votes in cases where the defending country is regarded as High Reputation Country Dothan (2014) . [^10] Countries with higher democracy index. See App K.
@@
Subjectivity refers to the disagreement due to personal opinions and values. We hypothesize that the GrandChamber, which consists of 17 judges, tends to more disagreement than the 7-judge Chamber.
@@ 5.3.1 Do proxy measures correlate with judges’ votes?
Table 1: Associations of proxy variables and entropy of the votes distribution ($\ast$: p < 0.05)
mic-F1 mac-F1 hm-F1
BERT 72.31 ± 4.09 65.53 ± 6.22 64.22 ± 3.46
LegalBERT 73.55 ± 0.62 68.75 ± 1.13 68.03 ± 1.29
Table 2: COC performance on test set
Tab 1 gives the results. It shows significantly lower agreement among judges’ votes for HighRepCountry and VulnApplicant. In other words, specific contexts of case facts correlate more with disagreement than Normative Uncertainty. Judges express disagreement related to how facts fit the norm because of circumstances related to the social standing of the applicant and the political situation of the defending State. This is an inherent characteristic of adjudication where judges qualify the context so as to fit the norm. The inclusion of a category related to the legal norm’s scope could have provided further insights.
@@
Interestingly, we found less disagreement among judges on KeyCase. Experts suggest this may stem from the Court’s assigning the importance score ex post, after interpretation and agreement are reached. The Court is more likely to designate a consensus case as KeyCase to maintain coherent jurisprudence, rather than a controversial case that could invite future applicants to base cases on disputed facts and dissenting opinions. Our DoF framework (App A) suggests that Case to Law uncertainty inherently perpetuates Law to Case uncertainty.
@@ 6 Judge-Model Misalignment
[…]
Table 3: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std ($\pm$) on test set. "/": COC finetuned models in § 6.1; "soft": models fine-tuned with soft-loss; "TS": models after Temperature Scaling. Results shown over 3 random seeds. See Tab 8 in App R for performance on dev set.
hm-F1 mean PVI t-value p-value
U. SV U. SV
B 66.19 ± 3.55 43.76 ± 4.23 0.94 0.06 3.12 0.002
L 69.3 ± 1.88 53.67 ± 4.59 0.77 -0.12 2.66 0.008
Table 4: COC prediction scores (hm-F1) and difficulty scores (mean PVI) for cases within the unanimous (U.) and SV categories. B: BERT; L: LegalBERT
6.1 Experimental Setup of COC
We assess the alignment of a model’s prediction difficulty [^12] Another approach would be assessing a model’s prediction uncertainty. Refer to App N for an explanation of our decision to abstain from using Uncertainty evaluation methods with judge disagreement, specifically COC models trained on ECtHR Task A|B. To address input length constraints, we employ a BERT variant of the hierarchical attention model [^13] See App O for model structure and implementation. Yang et al. (2016), as adapted from Santosh et al. 2022.
Models Santosh et al. 2022 encodes allegation information as a multi-hot vector concatenated with the text representation. By contrast, in this study, we textify allegation information as “Alleged Article: X”, where is X is a comma-separated list of articles and prepend to the input text. To explore the impact of domain-specific pre-training on uncertainty representations, we employ two BERT variants as the backbone in hierarchical models: (i) BERT Devlin et al. (2019) bert-base-uncased, trained on general English; (ii) LegalBERT Chalkidis et al. (2020) nlpaueb/legal-bert-base-uncased pretrained on English legal text, including ECtHR.
COC Performance: Tab 2 reports the results of classification performance. Notably, the model with legal-specific pre-training (LegalBERT) outperforms the one with general pre-training (BERT). Although the mic-F1 and mac-F1 performances closely align with those reported in Santosh et al. (2022); Xu et al. (2023b), differences of $\sim$ 2%, the hm-F1 performance demonstrate a noteworthy improvement of $\sim$10%. [^14] This can be partly attributed to our correction of labels (see § 4). Some variation may also be due to randomness in weight initialization. See our discussion in App M Importantly, hm-F1 exhibits particular sensitivity to allegation information. For the rest of this paper, we adopt hm-F1 as our primary metric for COC performance.
6.2 Difficulty of SV Cases in COC
Results: We present the average COC performance (hm-F1) and perceived difficulty (mean PVI scores) of the models in Tab 4. Both models exhibit lower hm-F1 and PVI scores for SV cases compared to unanimous cases. An independent t-test on the average PVI scores between the two groups reveals a significant difference ($p<0.05$). This indicates that SV cases are more challenging for the models than unanimous cases, aligning with human perceptions of difficulty. We also calculate the Pearson correlation coefficients between PVI scores and the entropy of judges’ vote distribution of SV cases. For BERT, the correlation coefficient (r-value) is -0.089, with a p-value of 0.38. LegalBERT has a r-value of -0.068 and a p-value of 0.48. The results reveal a negative correlation between PVI scores and the degree of disagreement among judges, consistent with our expectations. However, the observed correlation is very weak ($|r|<0.1$) and statistically insignificant ($p>0.05$). This suggests that, while models can capture differences in difficulty between unanimous and SV cases, they struggle to accurately represent the nuanced degree of judges’ disagreement. Therefore, we next evaluate the models’ confidence and soft-label training.
6.3 Calibration of SV Cases
@@ 6.3.1 Evaluation Metrics
Figure 3: The distribution over instance-based absolute errors between probabilities for class 1 of the model vs human vote distributions. See Fig 8 in App R for more figures comparing human uncertainty to model uncertainty
Most current NLP research focuses on Confidence Calibration Jiang et al. (2021); Desai and Durrett (2020): The model should be unsure when it does not know the answer. A model is considered well confidence-calibrated if its prediction confidence aligns with its predictive accuracy, commonly evaluated against the human majority class. Yet there is a growing interest in accounting for human label variants and/or pluralistic values Plank (2022); Sorensen et al. (2023). Therefore we extend our evaluation to include Human Calibration Baan et al. (2022): The model should be unsure when humans are unsure about the answer. We consider a model well human-calibrated if the categorical distribution of predicted class probabilities align well with the actual human vote distribution.
@@ 6.3.2 Calibration Methods
For confidence calibration, we employ Temperature Scaling (TS; Guo et al. 2017). This simple yet widely used post-hoc method uses a single temperature parameter $t$ to scale the output logits of a classifier. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. For human calibration, we adopt the approach from Peterson et al. 2019; Uma et al. 2020 by finetuning with a Soft Loss Function. [^15] See App P for details of the Soft Loss Function. During training, models are exposed to “soft labels” derived from the judges’ vote distributions, serving as target distributions in a cross-entropy loss function.
Expected Calibration Error (ECE) is the most often used metric for confidence calibration Naeini et al. (2015); Guo et al. (2017). A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true probabilities. Refer to the App Q for further details on ECE.

### 16 Feb 2024 v2

@@ 2 Related Work
COC has been referred to as Legal Judgment Prediction (LJP) in previous research Medvedeva et al. (2020); T.y.s.s et al. (2023). There exist numerous works involving corpora from ECtHR Aletras et al. (2016); Chalkidis et al. (2019); Tyss et al. (2023). All of these approaches mainly focus on the COC performance of models, which is commonly measured against the human majority class. Our SV-ECtHR dataset extends this work and contains the nuanced judges’ vote information. We proceed to use it to systematically investigate disagreement among judges and the alignment of perceived difficulty between models and judges.
@@
Calibration Existing calibration studies primarily focus on a classifier’s awareness of its predictive performance, commonly evaluated against the human majority class Guo et al. (2017); Desai and Durrett (2020). Baan et al. 2022 argue that calibrating against the human majority class is not meaningful in settings with inherent HLV. In our work, we assess the performance and human calibration of COC models in the context of split-vote, a naturally occurring HLV. To the best of our knowledge, this is the first systematic exploration of human calibration in the legal domain.
@@ 3 Legal and Linguistic Backgrounds
The Legal Decision Making Process at the ECtHR begins with the applicants lodging their accusation, alleging one or more violations of articles of the European Convention on Human Rights (ECHR). This process aligns with Task B (allegation prediction) in the widely used LexGLUE benchmark Chalkidis et al. (2022a). After receiving the complaint, the court undertakes a review of the case, aiming to determine whether a violation has indeed occurred, corresponding to Task A (violation prediction) in LexGLUE. Cases falling under the purview of well-established ECtHR case-law are directed to a three-judge Committee, while others may find themselves before a seven-judge Chamber, where decisions are reached through a majority vote. In some circumstances, the Grand Chamber, comprising 17 judges, adjudicates cases referred to it by request. When rendering a judgment, the Court typically examines only the specific articles alleged by the applicant. This procedure aligns with Task A|B (violation prediction given allegation; Santosh et al. 2022). Our study evaluates the uncertainty of COC models for Task A|B, mirroring the real legal process.
Sources of Disagreement in Legal Decisions Recent research has delved into sources of HLV and disagreement across various NLP tasks from a linguistic perspective. The classic framework “Triangle of Reference" Ogden and Richards (1923) has been widely adopted to categorize disagreement sources in classification tasks Aroyo and Welty (2015); Jiang and Marneffe (2022). This concept primarily addresses the relationship between linguistic symbols and the corresponding objects they represent. To study uncertainty sources in NLG tasks, Baan et al. 2023 extend it to the “Double Triangle of Language Production” catering to the complexities of language generation. However, the legal decision-making process introduces an additional layer of complexity. Legal scholars conceptualize the decision-making process in case-law as circular Ichim (2019). To better capture this nature, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4 in App A), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the continuous reproduction of the main normative model over time.
Figure 2: Taxonomy of disagreement sources among judges, with two adapted task-agnostic levels from Xu et al. 2023b and our expanded split-vote specific subcategories and proxy variables (in dashed box).
@@ 4 The SV-ECHR Dataset
Dataset Analysis We then augment the ECHR A|B dataset Santosh et al. (2022) with our collected vote information according to the document ID. We name the augmented dataset SV-ECHR. It consists of 11k case fact descriptions along with target label information about which convention articles have been alleged to be violated (task B), and which the court has eventually found to be violated (task A), and the judges vote distribution of each alleged article. The dataset is chronologically split into training (2001–2016), validation (2016–2017) and test set (2017-2019) with 9k, 1k, 1k cases respectively. The label set includes 10 prominent ECHR convention articles. On average, each case has around 1.6 alleged articles. Among all 17,604 alleged case-alleged article pairs (hereafter pair), only round 7% are split-voted. These numbers underscore the significant label imbalance within the dataset. Additional statistics can be found in Tab 4 in App D. To investigate the extent of judges’ disagreement in SV-ECHR, we assess the entropies of vote distributions for each pair. For the detailed calculation and histogram of the entropy distribution, see App F. We see a large share ($\sim$ 60% of chamber votes) of single dissenting votes, where only one single judge voted differently than the six judge majority. A similar pattern is observed for Grand Chamber cases involving 17 judges. This pattern aligns with other observations made in empirical legal scholarship (Fobbe 2022, see App G).
@@ 5.1.1 The Data
Genuine Ambiguity is attributed to Normative Uncertainty of case law in the context of legal NLP (Xu et al. 2023b), which emerges when the court is presented with the possibility of justifying an outcome through multiple legal source interpretations and argumentation. Its occurrence is not uncommon in ECtHR judgments due to the deliberate drafting of the convention in a ‘flexible’ manner to allow tailoring the interpretation to domestic specificities based on the subsidiarity and margin of appreciation principles. It should be noted that there exists an analog factual ambiguity where the facts of the case are unclear based on limited (or contradicting) evidence provided. As the ECtHR does not engage in evidentiary reasoning, this aspect is out of scope for this work.
@@ 5.1.2 The Annotator/Judge
A comparable liberal vs conservative investigation in the ECtHR context would not be useful. It is composed of 46 Member States, each with its own spectrum of political ideologies, more pluralist and complicated than the liberal-conservative divide. Moreover, there is no temporal stability in the judges’ behavior because they have a limited mandate and act in different bench formations (chamber, grand chamber, etc.). Focusing on individual ideologies to the exclusion of variables accounting for the wider court dynamics could not support accurate analysis.
@@ 5.3 Proxies & Disagreement Correlation
HighRep (High Reputation Country): Previous legal and political science work indicates ECtHR Judges tend to have more split votes in cases where the defending country is regarded as High Reputation Country [^10] Countries with higher democracy index. See App K. Dothan (2014) .
@@
Subjectivity refers to the disagreement due to personal opinions and values. We propose that the 17-judge Grand Chamber is more prone to disagreement than the 7-judge chamber. This is due to the inherent pluralism of human values in larger groups, which naturally fosters a broader range of viewpoints and thus, increased subjectivity. As explained in sec. 5.1.2, modeling specific ideological dimensions of individual judges is out of scope for this work. We account for some limited political dynamics by including the democratic score of the state against which the claim is brought (the HighRep variable above).
@@ 5.3.1 Do proxy measures correlate with judges’ votes?
Table 1: Associations between proxy variables and vote distribution entropy. *: p < 0.05. Green highlighted: confirming our hypothesis. Red highlighted: contradicting our hypothesis.
Tab 1 shows significantly lower agreement among judges’ votes for HighRep and VulnApplicant (p-value < 0.05). In other words, specific contexts of case facts correlate more with disagreement than Normative Uncertainty. Judges express disagreement related to how facts fit the norm because of circumstances related to the social standing of the applicant and the political situation of the defending State. This is an inherent characteristic of adjudication where judges qualify the context so as to fit the norm. The inclusion of a category related to the legal norm’s scope could have provided further insights.
@@
Interestingly, we found less disagreement among judges on KeyCase. Experts suggest this may stem from the Court’s assigning the importance score ex post, after interpretation and agreement are reached. The Court is more likely to designate a consensus case as KeyCase to maintain coherent jurisprudence, rather than a controversial case that could invite future applicants to base cases on disputed facts and dissenting opinions. As reflected in our DoF framework (App A), Case to Law uncertainty inherently perpetuates Law to Case uncertainty.
@@ 6 Judge-Model Misalignment
6.1 Experimental Setup of COC
We assess the alignment of a model’s prediction difficulty [^12] Another approach would be assessing a model’s prediction uncertainty. Refer to App O for an explanation of our decision to abstain from using Uncertainty evaluation methods with judge disagreement, specifically COC models trained on ECtHR Task A|B. To address input length constraints, we employ a LegalBERT Chalkidis et al. (2020) variant of the hierarchical attention model [^13] See App P for model structure and implementation. Yang et al. (2016), as adapted from Santosh et al. 2022.
Models Santosh et al. 2022 encodes allegation information as a multi-hot vector concatenated with the text representation. By contrast, in this study, we textify allegation information as “Alleged Article: X”, where is X is a comma-separated list of articles and prepend to the input text. We employ the LegalBERT nlpaueb/legal-bert-base-uncased as the backbone in hierarchical models [^14] To explore the impact of domain-specific pre-training, we did also experiment on BERT. See App M..
COC Performance We run experiments across three random seeds. The hierarchical LegalBERT achieves 73.55 ± 0.62 for mic-F1, 68.75 ± 1.13 for mac-F1 and 68.03 ± 1.29 for hm-F1. Although the mic-F1 and mac-F1 performances closely align with those reported in Santosh et al. (2022); Xu et al. (2023b), differences of $\sim$ 2%, the hm-F1 performance demonstrate a noteworthy improvement of $\sim$10%. [^15] This can be partly attributed to our correction of labels (see § 4). Some variation may also be due to randomness in weight initialization. See our discussion in App N Importantly, hm-F1 exhibits particular sensitivity to allegation information. For the rest of this paper, we adopt hm-F1 as our primary metric for COC performance.
6.2 Difficulty of SV Cases in COC
hm-F1 mean PVI t-value p-value
U. SV U. SV
LegalBert 69.3 ± 1.88 53.67 ± 4.59 0.77 -0.12 2.66 0.008
Table 2: COC prediction (hm-F1) and difficulty scores (mean PVI) for unanimous (U.) and SV cases.
Results: Tab 2 presents the average COC performance (hm-F1) and model-perceived difficulty (mean PVI scores). The model exhibits lower hm-F1 and PVI scores for SV cases compared to unanimous cases. An independent t-test on the average PVI scores between the two groups reveals a significant difference ($p<0.05$). This indicates that SV cases are more challenging for the models than unanimous cases, aligning with human perceptions of difficulty. We also calculate the Pearson correlation coefficients between PVI scores and the entropy of judges’ vote distribution of SV cases. LegalBERT has a correlation coefficient (r-value) of -0.068 and a p-value of 0.48. The results reveal a negative correlation between PVI scores and the degree of disagreement among judges, consistent with our expectations. However, the observed correlation is very weak ($|r|<0.1$) and statistically insignificant ($p>0.05$). This suggests that, while the model can capture differences in difficulty between unanimous and SV cases, it struggle to accurately represent the nuanced degree of judges’ disagreement. Therefore, we next evaluate the models’ confidence and soft-label training.
6.3 Calibration of SV Cases
@@
Table 3: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std ($\pm$) on test set. "/": COC finetuned models in § 6.1; "soft": models fine-tuned with soft-loss; "TS": models after Temperature Scaling. Results shown over 3 random seeds. See Tab 8 in App S for performance on dev set.
@@ 6.3.1 Evaluation Metrics
Figure 3: The distributions over probabilities for class 1 of the models vs human vote distributions (row 1) and distCE (row 2). See Fig 8 in App S for more figures comparing human uncertainty to model uncertainty
Most current NLP research focuses on Confidence Calibration Jiang et al. (2021); Desai and Durrett (2020): The model should be unsure when it does not know the answer. A model is considered well confidence-calibrated if its prediction confidence aligns with its predictive accuracy, commonly evaluated against the human majority class. Yet there is a growing interest in accounting for HLV and/or pluralistic values Plank (2022); Sorensen et al. (2023). Therefore we extend our evaluation to include Human Calibration Baan et al. (2022): The model should be unsure when humans are unsure about the answer. We consider a model well human-calibrated if the categorical distribution of predicted class probabilities align well with the actual human vote distribution.
@@ 6.3.2 Calibration Methods
For confidence calibration, we employ Temperature Scaling (TS; Guo et al. 2017). This simple yet widely used post-hoc method uses a single temperature parameter $t$ to scale the output logits of a classifier. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. For human calibration, we adopt the approach from Peterson et al. 2019; Uma et al. 2020 by finetuning with a Soft Loss Function. [^16] See App Q for details of the Soft Loss Function. During training, models are exposed to “soft labels” derived from the judges’ vote distributions, serving as target distributions in a cross-entropy loss function.
Expected Calibration Error (ECE) is the most often used metric for confidence calibration Naeini et al. (2015); Guo et al. (2017). A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true probabilities. Refer to the App R for further details on ECE.